# Dose-aware Diffusion Model for 3D Low-dose PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v1

@@ V Ablation studies
TABLE III: Quantitative assessment for different ablation studies. The best results among different low-count levels are marked in red. The second best results are marked in blue. Black voxels were removed for calculations.
@@
TABLE IV: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations.
@@ Acknowledgment
All authors declare that they have no known conflicts of interest in terms of competing financial interests or personal relationships that could have an influence or are relevant to the work reported in this paper.

### v2

@@ III-E Albation studies
TABLE III: Quantitative assessment for different ablation studies. The best results among different low-count levels are marked in red. The second best results are marked in blue. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
@@
TABLE IV: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
@@
IV Discussion and Conclusion
We introduced DDPET-3D, a dose-aware diffusion model for 3D low-dose PET imaging. Evaluated on a large-scale, multi-center, and cross-scanner dataset, low-dose/low-count denoised images generated by DDPET-3D performed favorably or comparably relative to the images reconstructed with 100% full-count. Tested on images acquired using three state-of-the-art PET/CT systems from different vendors and confirmed with reader studies, DDPET-3D consistently produced superior denoised results. Tested on images acquired at a different hospital, DDPET-3D also demonstrated its potential to be easily transferred to another medical center without further network fine-tuning. In previous literature, about 77% of deep-learning-based healthcare papers are limited to institutional or regionally distributed dataset [35]. Cross-center generalizability would be crucial to implement deep-learning algorithms in clinical practice.
Compared to previous diffusion models which are time-consuming to reconstruct the entire 3D volumes, once the DDPET-3D is fully-trained, the denoising process is efficient and can be finished within 15 min using one NVIDIA A40 GPU.
Confirmed with reader studies by professional nuclear physicians, the proposed DDPET-3D can provide a similar or better image quality compared with normal-dose images in terms of overall subjective visual impression. DDPET-3D effectively suppressed background noise and false hot spots in certain patients, making true lesions/hot spots easier to identify.
[…]
In conclusion, the proposed DDPET-3D provides better or similar image quality in different low-dose settings compared to images acquired in a normal-dose setting as measured by an expert visual preference study. The experimental results showed that DDPET-3D achieved superior performance compared with previous methods. While this study focuses on PET image denoising, we believe that the proposed method could be easily extended to other 3D reconstruction tasks for different imaging modalities.
@@ Acknowledgment
Ramsey D. Badawa, Benjamin A. Spencer, and Reimund Bayerlein declared that UC Davis has a revenue sharing agreement with United Imaging Healthcare.
Other authors declare that they have no known conflicts of interest in terms of competing financial interests or personal relationships that could have an influence or are relevant to the work reported in this paper.